k/15+5=k/6-10 solve for k
Answer by checkley77(12844)

k/15+5=k/6-10
K/15-K/6=-10-5
(2K-5K)/30=-15
-3K/30=-15
-3K=-15*30
-3K=-450
K=-450/-3
K=150 ANS.
PROOF:
150/15+5=150/6-10
10+5=25-10
15=15